The slope of f(x)=x3−12x2+36x−48 is zero at which value/s of x?


1
Expert's answer
2022-05-17T08:16:47-0400
"y'=(x^3-12x^2+36x-48)'"

"=3x^2-12(2x)+36(1)-0"

"=3x^2-24x+36"


If the slope is zero 


"3x^2-24x+36=0"

"x^2-8x+12=0"

"(x-2)(x-6)=0"

"x_1=2, x_2=6"

The slope of "f(x)=x^3\u221212x^2+36x\u221248" is zero at "x=2" or at "x=6."